A charge q is placed at the centre of the open end of a cylindrical vessel as shown in fig. The flux of the
electric field through the surface of the vessel is -

Grade:

2 Answers

Sunil Kumar FP
9 years ago
A charge q is placed at the centre of the open end of a cylindrical vessel as shown in fig. The flux of the electric field through the surface of the vessel is q/2\xi _{0}
SANJAY SINGHANIA
7 years ago
A charge q is placed at the centre of the open end of cyclinder then charge inclose by cyclinder is half of the charge E=q/2/€o=q/2€o